Use the molar bond enthalpy data in the table to e
Answer #1
Concepts and reason

This problem can be solved by using the general equation for the heat of a reaction.

Fundamentals

The heat of a reaction or change in enthalpy of a reaction can be estimated from bond energy. Heat energy is estimated after balancing the given equation.

Bond energy is defined as the amount of energy required to break the bond of 1 mole of a gaseous compound to produce its constituents.

The change in enthalpy of a reaction is the difference of summation of bond energies of bond broken and the summation of the bond energies of the bonds formed. It is written as follow.

ΔHo=ΔH(reactantbondsbroken)ΔH(productbondsformed){\rm{\Delta }}{H^o} = \sum {{\rm{\Delta }}{H_{\left( {{\rm{reactant bonds broken}}} \right)}}} - \sum {{\rm{\Delta }}{H_{\left( {{\rm{product bonds formed}}} \right)}}} …… (1)

The reaction is as follows:

CCl4(g)+2F2(g)CF4(g)+2Cl2(g){\rm{CC}}{{\rm{l}}_{\rm{4}}}\left( g \right) + 2{{\rm{F}}_2}\left( g \right) \to {\rm{C}}{{\rm{F}}_4}\left( g \right) + 2{\rm{C}}{{\rm{l}}_{\rm{2}}}\left( g \right)

First calculate the bond energy of the reactants or the bonds broken as follow.

ΔH(reactantbondsbroken)=4ΔH(CCl)+2ΔH(FF)\sum {{\rm{\Delta }}{H_{\left( {{\rm{reactant bonds broken}}} \right)}}} = 4{\rm{\Delta }}H\left( {{\rm{C}} - {\rm{Cl}}} \right) + 2{\rm{\Delta }}H\left( {{\rm{F}} - {\rm{F}}} \right)

Substitute 331kJmol1{\rm{331 kJ mo}}{{\rm{l}}^{ - 1}} for ΔH(CCl){\rm{\Delta }}H\left( {{\rm{C}} - {\rm{Cl}}} \right) and 155kJmol1{\rm{155 kJ mo}}{{\rm{l}}^{ - 1}} for ΔH(FF){\rm{\Delta }}H\left( {{\rm{F}} - {\rm{F}}} \right) .

ΔH(reactantbondsbroken)=4(331kJmol1)+2(155kJmol1)=1634kJmol1\begin{array}{c}\\\sum {{\rm{\Delta }}{H_{\left( {{\rm{reactant bonds broken}}} \right)}}} = 4\left( {{\rm{331 kJ mo}}{{\rm{l}}^{ - 1}}} \right) + 2\left( {{\rm{155 kJ mo}}{{\rm{l}}^{ - 1}}} \right)\\\\ = 1634{\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\end{array}

Now, calculate bond energy of product bonds formed.

ΔH(productbondsformed)=4ΔH(CF)+2ΔH(ClCl)\sum {{\rm{\Delta }}{H_{\left( {{\rm{product bonds formed}}} \right)}}} = {\rm{4\Delta }}H\left( {{\rm{C}} - {\rm{F}}} \right) + 2{\rm{\Delta }}H\left( {{\rm{Cl}} - {\rm{Cl}}} \right)

Substitute 439kJmol1{\rm{439 kJ mo}}{{\rm{l}}^{ - 1}} for ΔH(CF){\rm{\Delta }}H\left( {{\rm{C}} - {\rm{F}}} \right) and 243kJmol1{\rm{243 kJ mo}}{{\rm{l}}^{ - 1}} for ΔH(ClCl){\rm{\Delta }}H\left( {{\rm{Cl}} - {\rm{Cl}}} \right) .

ΔH(productbondsformed)=4(439kJmol1)+2(243kJmol1)=2242kJmol1\begin{array}{c}\\\sum {{\rm{\Delta }}{H_{\left( {{\rm{product bonds formed}}} \right)}}} = 4\left( {{\rm{439 kJ mo}}{{\rm{l}}^{ - 1}}} \right) + 2\left( {{\rm{243 kJ mo}}{{\rm{l}}^{ - 1}}} \right)\\\\ = 2242{\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\end{array}

Now, substitute 1634 kJ for ΔH(reactantbondsbroken)\sum {{\rm{\Delta }}{H_{\left( {{\rm{reactant bonds broken}}} \right)}}} and 2242 kJ for ΔH(productbondsformed)\sum {{\rm{\Delta }}{H_{\left( {{\rm{product bonds formed}}} \right)}}} in equation (1).

ΔHo=ΔH(reactantbondsbroken)ΔH(productbondsformed)=(16342242)kJmol1=608kJmol1\begin{array}{c}\\{\rm{\Delta }}{H^o} = \sum {{\rm{\Delta }}{H_{\left( {{\rm{reactant bonds broken}}} \right)}}} - \sum {{\rm{\Delta }}{H_{\left( {{\rm{product bonds formed}}} \right)}}} \\\\ = \left( {1634 - 2242} \right){\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\\ = - 608{\rm{ kJ mo}}{{\rm{l}}^{ - 1}}\\\end{array}

Ans:

The heat of the reaction is 608kJmol1 - 608{\rm{ kJ mo}}{{\rm{l}}^{ - 1}} .